If the numbers 70 and 105 are both multiples of a prime number n, how many possible values are there for n?

Answer by Edwin McCravy(20086)

 70 = 2×35 = 2×5×7

105 = 3×35 = 3×5×7

They are both multiples of the 2 prime numbers 5 and 7.

Answer: There are 2 possible values for n.

Edwin